Solve Problems Using Linear Systems
16. One type of fertilizer has 30% nitrogen and a second type has 15% nitrogen. If a farmer needs 600kg of fertilizer that is 20% nitrogen, how much of each type should the farmer mix together?
Thank you very much.

Answer by checkley77(12844)

.30X+.15(600-X)=600*.20
.30X+90-.15X=120
.15X=120-90
.15X=30
X=30/.15
X=2000 KG. OF 30% NITROGEN.
600-200=400 KG. OF .15% NITROGEN.
PROOF:
.30*200+.15*400=120
60+60=120